### v1

@@ 0.C.3 Computing a Collection with tbn(G) Bends
After this procedure, $G^{1}_{C}$ and $G^{2}_{C}$ have no bad 3-legged cycle as, for every demanding cycle, we placed a dummy either in one of its interesting contour paths or in the edge individuated by Lemma 16 in both $G^{1}_{C}$ and $G^{2}_{C}$. We placed $2|D(G_{C})|+count_{k}$ dummies. We increment $count_{i}$ with respect to $count_{i-1}$ ($i\in[1,k]$) only when $\beta(C^{\prime})=\emptyset$. By 1 of the enriched genealogical tree $T^{+}(C)$, we placed exactly $2|D(G_{C})|+|D_{\exp}(G_{C})|$ subdivision vertices.
We now show that the placement satisfies (ii). We consider the cycles that are contained in $D_{\operatorname{short}}(G_{C})$. Observe that, each $C^{*}\in D_{\operatorname{short}}(G_{C})$ has at last two edges belonging to interesting contour paths. Also, only one edge $e^{*}$ lies along $P$. Hence, according to the non-expensive edge-selection rule, we subdivided such edge $e^{*}$ in exactly one between $G^{1}_{C}$ or $G^{2}_{C}$. We now have that every cycle in $D_{\operatorname{short}}(G_{C})$ has exactly one dummy contained in $P$ in the collection. Hence, at most $2t-|D_{\operatorname{short}}(G_{C})|$ dummies are incident to $f_{\operatorname{ext}}$. We show now that exactly such number of dummies are incident to $f_{\operatorname{ext}}$. Namely, it remains to show that the other exposed interesting contour paths of non-short demanding cycles always imply a subdivision vertex on $f_{\operatorname{ext}}$. Consider critical cycles: in this case, in each one of $G^{1}_{C}$ and $G^{2}_{C}$, we either subdivided an exposed contour path contained in $P$ or (if $C^{\prime}$ is expensive) we subdivided one of its legs incident to $f_{ext}$ following the property of Lemma 16. For the demanding cycles with two edges contained in an interesting contour path that are not short and with an exposed contour path, we subdivided an edge incident to $f_{\operatorname{ext}}$ in both $G^{1}_{C}$ and $G^{2}_{C}$ by the non-expensive edge-selection rule. Hence, the collection has exactly $2t-|D_{\operatorname{short}}(G_{C})|$ edges placed on the exposed contour path $P$ of $C$.
Part 2. Here we show that the dummies placed above are sufficient for $G^{1}_{C}$ and $G^{2}_{C}$ to not to have bad cycles, the remaining of which are the 0-legged cycle that is the external face of $G_{C}$ and the 2-legged cycles caused by the presence of dummies. We prove it for $G^{1}_{C}$, the proof for $G^{2}_{C}$ is the same. Observe that $G_{C}$ is a triconnected cubic graph with three different edges of the external face that are subdivided. These vertices are the ones incident to the legs of $C$ in $G$ and here we can consider them like dummies that we denote by $d_{1}$, $d_{2}$, and $d_{3}$.
@@
In the procedure of Part 1 of the proof we subdivided edges that were contained in interesting contour paths, i.e., contour paths of bad cycles of 3-legged cycles. These contour paths had no edge subdivided and consequently we did not subdivide any edge having one between $d_{1}$, $d_{2}$, or $d_{3}$. Hence, there are at least 3 edges subdivided in $G^{1}_{C}$ and by Lemma 6, there are no bad 2-legged cycles.
We now prove that there are no pairwise intersecting demanding cycle contained in $G_{C}$. Suppose by contradiction that it is the case. Then, we slightly modify graph $G$ so that we have a contradiction to Condition (3) of Lemma 7. We proceed as follow: we consider any edge not contained in $G_{C}$ and incident to $f_{\operatorname{ext}}$ and we subdivide it twice inserting two vertices $v_{1}$ and $v_{2}$; we subdivide any other edge incident to the same internal face with a vertex $v_{3}$; We add a vertex $v_{4}$ and the edges $v_{1}v_{4}$, $v_{2}v_{4}$, and $v_{3}v_{4}$ so that the graph we obtain is plane. The obtained graph is also triconnected cubic and it has a 3-legged cycle formed by three edges $\{v_{1}v_{2},v_{1}v_{4},v_{2}v_{4}\}$ incident to the external face of $G$ and not contained in $G_{C}$. Such 3-legged cycle is a demanding cycle, since no vertex is contained in its interior. Since $I(G)\not=\emptyset$ by hypothesis, we have a contradiction to Condition (3) of Lemma 7.
@@
See 4.1
Proof
We prove that we can compute in $O(n)$ time a collection with at most $q=2|D(G)|+|D_{\exp}(G)|+8-\min\{8,2|D_{\operatorname{ext}}(G)|-|D_{\operatorname{short}}(G)|\}$ bends. Our strategy is first creating two copies of $G$, $G^{1}$ and $G^{2}$, and then placing on them overall exactly $q$ dummies. We prove that none of the two graphs has a bad cycle; then, by Lemma 4 and since $q\leq\operatorname{tbn}(G)$ by Lemma 5, this directly gives us an unbent collection with $\operatorname{tbn}(G)$ bends. Let $f_{\operatorname{ext}}$ be the external face of $G$. We first prove the theorem for a simple case, where there is no 3-legged cycle with edges incident to $f_{\operatorname{ext}}$ (hence, in particular, $D_{\operatorname{ext}}(G)=\emptyset$) and $f_{\operatorname{ext}}$ is incident to at least four edges. If these conditions hold, we say that $G$ is nice.
G is nice. Let $G^{1}$ and $G^{2}$ be two copies of $G$. Observe that in this case, since $G$ is nice, $D_{\operatorname{ext}}(G)=\emptyset$ and consequently $02|D_{\operatorname{ext}}(G)|-|D_{\operatorname{short}}|=$ 0. Hence, by Lemma 5, $\operatorname{tbn}\geq q=2|D(G)|+|D_{\exp}(G)|+8$. We show how to place $q$ dummies overall so that $G^{1}$ and $G^{2}$ have no bad cycle.
We compute the maximal 3-legged cycles of $G$ by Lemma 12. Let $C$ be one of such maximal 3-legged cycles. Consider the subgraphs $G^{1}_{C}$ and $G^{2}_{C}$ of $G^{1}$ and $G^{2}$, respectively. We place dummies along the edges of $G^{1}_{C}$ and $G^{2}_{C}$ by the algorithm whose existence is stated by Lemma 17. We perform this procedure for every maximal 3-legged cycle. We have that since $G$ is nice and by Lemma 7(3), the maximal 3-legged cycles of $G$ are disjoint. Hence, by Lemma 17, we placed exactly $q^{\prime}=2|D(G)|+|D_{\exp}(G)|$ dummies along the edges of $G^{1}$ and of $G^{2}$. Consider the external face of $G$ and select four edges $e_{1}^{1},e_{2}^{1},e_{1}^{2},e_{2}^{2}$ incident to $f_{\operatorname{ext}}$ such that both the pair $e_{1}^{1},e_{2}^{1}$ and the pair $e_{1}^{2},e_{2}^{2}$ are independent pairs of edges (this is always possible, as $G$ is nice and consequently it has at least four edges incident to the external face). We place two dummies along $e_{1}^{1}$ and $e_{2}^{1}$ in $G^{1}$ and two dummies along $e_{2}^{1}$ and $e_{2}^{2}$ in $G^{2}$. The external cycle of both $G^{1}$ and $G^{2}$ is not a bad cycle, as we just placed four dummies in each of them. Also, there is no other bad cycle by Lemma 6. Observe that this procedure places $q=2|D(G)|+|D_{\exp}(G)|+8$ dummies. The linear time complexity follows from Lemma 4, by the fact that when $G$ is nice maximal 3-legged cycles of $G$ are independent (as observed before), and by Lemma 17.
@@
We first observe that these cases are mutually exclusive. The fact that (i) and (iii) are mutually exclusive directly comes from Lemma 7(3). Concerning (ii) and the other two, it suffices to prove that when the external cycle is a 3-cycle there is no 3-legged incident to $f_{\operatorname{ext}}$. Suppose by contradiction that $f_{\operatorname{ext}}$ is incident to 3 edges and that there is a 3-legged $C$ having $f_{\operatorname{ext}}$ as leg face. Then two of the edges incident to $f_{\operatorname{ext}}$ are legs of $C$ and at least one of them is an edge of $C$ forming one of its contour paths. Since these are 3 edges already, incident to $f_{\operatorname{ext}}$, it means that all the 3 legs of $C$ are incident to a same vertex. A contradiction (see the definition of $k$-legged cycle).
(i) Intersecting Case. Refer to Fig. 17, where $C_{1}$, $C_{2}$ and $C_{3}$ are intersecting demanding cycles and $C_{1}^{t}$, $C_{2}^{t}$, and $C_{3}^{t}$ are their respective twins. We compute all the intersecting demanding cycles by Lemma 12, computing all the demanding and subtracting from this set the demandings in $D(G)$. We select two intersecting demanding cycles $C_{1}$ and $C_{2}$ and the relative twins $C_{1}^{t}$ and $C_{2}^{t}$. We then consider $C_{1}$ and $C_{1}^{t}$, we select two edges $e\in C_{1}\cap f_{\operatorname{ext}}$ and $e^{\prime}\in C_{1}^{t}\cap f_{\operatorname{ext}}$. We place in $G^{1}$ two dummies in $e$ and two dummies in $e^{\prime}$. By Lemma 7(2) there is no 3-legged bad cycle that is intersecting in $G^{1}$. Also the external face is now not a bad cycle in $G^{1}$ and by Lemma 6 there is no bad 2-legged cycle. We do the same considering $C_{2}$, $C_{2}^{t}$, and $G^{2}$. Concerning the internal 3-legged cycles, we handle them with the nice-placement. By Lemma 7(3) we have $D_{\operatorname{ext}}(G)=\emptyset$. Since $D_{\operatorname{ext}}(G)=\emptyset$, as in the case when $G$ is nice, we have $\operatorname{tbn}(G)\geq q=2|D(G)|+|D_{\exp}(G)|+8$ by Lemma 5. We have that with this placement strategy we placed $q^{\prime}=2|D(G)|+|D_{\exp}(G)|$ internal dummies by Lemma 17 as in the case when $G$ is nice and, considering the $8$ we placed in the external face to handle the intersecting demanding cycles, we placed overall $q=q^{\prime}+8=2|D(G)|+|D_{\exp}(G)|+8$. Since we have no bad cycle, as we handled the intersecting applying Lemma 7(2) and since the 3-legged are handled by the same procedure used when $G$ is nice, the lemma follows. Concerning the computational time, the fact that this procedure takes linear time comes from Lemma 12 and since the nice-placement takes linear time (as proved in the case when $G$ is a nice graph).
@@
(ii) 3-Cycle Case. As observed while showing that the three cases we are considering are mutually exclusive, we have that in this case there is no 3-legged cycle having $f_{\operatorname{ext}}$ as leg face and consequently no intersecting cycle and $D_{\operatorname{ext}}(G)=\emptyset$. We now have 3 edges in the external face and 4 dummies to place so that it is not a bad cycle, but here, the situation changes with respect to the case when $G$ is nice, since we cannot directly use Lemma 6. In fact, since the external face has 3 edges we cannot place the 4 dummies in 3 different edges or in at least 2 edges with 2 dummies each that are independent, as described by Lemma 6. Hence, here we do not use directly Lemma 5, but we rely on the fact that the nice-placement places dummies optimally in $G^{1}\setminus e$ and $G^{2}\setminus e$ and on the fact that the placement of dummies on edges of $f_{\operatorname{ext}}$ has bounded number of possibilities. Namely, there are three possibilities:
@@
We first show how to handle the 1-2-placement. Consider first $G^{1}$. Refer to Fig. 18(a). We subdivide $e=uv$ with four dummies $d_{1},...,d_{4}$ so that $d_{1}$ is adjacent to $u$ and $d_{4}$ is adjacent to $v$. Let $C_{\operatorname{bad}}$ be the external cycle of $G^{1}\setminus e$. We have that $C_{\operatorname{bad}}$ is a bad 2-legged cycle and we need to place two dummies on it. We are going to show how to subdivide $C_{\operatorname{bad}}$ optimally. Observe that $C_{\operatorname{bad}}$ has two legs, which are $(u,d_{1})$ and $(v,d_{4})$. Suppose that $e^{\prime}$ and $e^{*}$, which are the two edges different from $e$ incident to $f_{\operatorname{ext}}$, are incident to $u$ and $v$ respectively. Let $e^{\prime}_{\mathrm{int}}$ and $e^{*}_{\mathrm{int}}$ be the other edges incident to $u$ and $v$, which are internal edges of $G$ Let $\Pi$ be a path such that $C_{\operatorname{bad}}=e^{\prime}\cup e^{*}\cup e^{*}_{\mathrm{int}}\cup e^{\prime}_{\mathrm{int}}\cup\Pi$ and let $f_{e}$ be the face incident to $e$ and internal in $G$.
We now use the nice-placement in $G\setminus e$. We just placed $q^{\prime}=2|D(G)|+|D_{\exp}(G)|$. By Lemma 11 the interesting contour paths and, consequently, the legs of the demanding cycles incident to $f_{e}$ lie along $\Pi$. Hence, if such demanding cycles exists, nice-placement subdivided $\Pi$ already by Lemma 17(2). If $C_{\operatorname{bad}}$ has 2 dummies in $\Pi$, we do nothing, and the cost is optimum as we did not add dummies with respect to the nice-placement. Otherwise, suppose that that $C_{\operatorname{bad}}$ does not have 2 dummies.
@@
We now consider $G^{2}$. Refer to Fig. 18(b). Let $w$ be the vertex of $G$ incident to $e^{\prime}$ and $e^{*}$. We have to consider the degenerate 3-legged cycle $C_{\operatorname{bad}}^{\prime}$ of $G$ whose three legs are incident to $w$, since two of its legs are incident to dummies. In this case we have to place just one dummy $d^{\prime}$. As before, we can observe that either the nice-placement placed such dummy in one of its contour paths $\Pi^{\prime}$ and $\Pi^{*}$ (sharing a face with $e^{\prime}$ and $e^{*}$, respectively, and defined similarly as $\Pi$) previously, or we can place it in any edge of such contour paths. We have that we did not place a dummy in a same edge in $G^{1}$ and $G^{2}$. In fact, $\Pi^{\prime}$ and $\Pi^{*}$ are disjoint from $\Pi$.
@@
(iii) External-Demanding Case: $D_{\operatorname{ext}}(G)\not=\emptyset$. As there is no intersecting cycle, since the cases we are considering are mutually exclusive, we have that the maximal 3-legged cycles are independent. Hence, we can use the nice-placement so that now every 3-legged cycle is not a bad cycle. We placed exactly $q^{\prime}=2|D(G)|+|D_{\exp}(G)|$ dummies. We have now to see what happens in the external face $f_{\operatorname{ext}}$. By Lemma 17(2) we placed a dummy in each of $G^{1}$ $G^{2}$ for every cycle in $D_{\operatorname{ext}}(G)\setminus D_{\operatorname{short}}(G)$ and just one dummy (either in $G^{1}$ or $G^{2}$) for the cycles in $D_{\operatorname{short}}(G)$. The remaining dummies so that the external face has at least four vertices can be placed in other edges in $G^{1}$ and $G^{2}$. We have that by Lemma 17 and by Lemma 5 we placed $\operatorname{tbn}(G)$ dummies. Also, by Lemma 7 and by Lemma 17 there is no bad cycle.
@@
Corollary 2
Let $G$ be a planar triconnected $3$-graph with $n$ vertices. There exists an $O(n^{2})$-time algorithm that computes an unbent collection of $G$ with $\operatorname{tbn}(G)$ bends. Also, the collection has size two which is optimal.

### v2

@@ C.3 Computing a Collection with tbn(G) Bends
After this procedure, $G^{1}_{C}$ and $G^{2}_{C}$ have no bad 3-legged cycle as, for every demanding cycle, we placed a dummy either in one of its interesting contour paths or in the edge individuated by Lemma C.16 in both $G^{1}_{C}$ and $G^{2}_{C}$. We placed $2|D(G_{C})|+count_{k}$ dummies. We increment $count_{i}$ with respect to $count_{i-1}$ ($i\in[1,k]$) only when $\beta(C^{\prime})=\emptyset$. By 1 of the enriched genealogical tree $T^{+}(C)$, we placed exactly $2|D(G_{C})|+|D_{\exp}(G_{C})|$ subdivision vertices.
We now show that the placement satisfies (ii). We consider the cycles that are contained in $D_{\operatorname{short}}(G_{C})$. Observe that, each $C^{*}\in D_{\operatorname{short}}(G_{C})$ has at last two edges belonging to interesting contour paths. Also, only one edge $e^{*}$ lies along $P$. Hence, according to the non-expensive edge-selection rule, we subdivided such edge $e^{*}$ in exactly one between $G^{1}_{C}$ or $G^{2}_{C}$. We now have that every cycle in $D_{\operatorname{short}}(G_{C})$ has exactly one dummy contained in $P$ in the collection. Hence, at most $2t-|D_{\operatorname{short}}(G_{C})|$ dummies are incident to $f_{\operatorname{ext}}$. We show now that exactly such number of dummies are incident to $f_{\operatorname{ext}}$. Namely, it remains to show that the other exposed interesting contour paths of non-short demanding cycles always imply a subdivision vertex on $f_{\operatorname{ext}}$. Consider critical cycles: in this case, in each one of $G^{1}_{C}$ and $G^{2}_{C}$, we either subdivided an exposed contour path contained in $P$ or (if $C^{\prime}$ is expensive) we subdivided one of its legs incident to $f_{ext}$ following the property of Lemma C.16. For the demanding cycles with two edges contained in an interesting contour path that are not short and with an exposed contour path, we subdivided an edge incident to $f_{\operatorname{ext}}$ in both $G^{1}_{C}$ and $G^{2}_{C}$ by the non-expensive edge-selection rule. Hence, the collection has exactly $2t-|D_{\operatorname{short}}(G_{C})|$ edges placed on the exposed contour path $P$ of $C$.
Part 2. Here we show that the dummies placed above are sufficient for $G^{1}_{C}$ and $G^{2}_{C}$ to not to have bad cycles, the remaining of which are the 0-legged cycle that is the external face of $G_{C}$ and the 2-legged cycles caused by the presence of dummies. We prove it for $G^{1}_{C}$, the proof for $G^{2}_{C}$ is the same. Observe that $G_{C}$ is a plane triconnected cubic graph with three different edges of the external face that are subdivided. These vertices are the ones incident to the legs of $C$ in $G$ and here we can consider them as dummies; let $d_{1}$, $d_{2}$, and $d_{3}$ denote them.
@@
In the procedure of Part 1 of the proof we subdivided edges that were contained in interesting contour paths, i.e., contour paths of bad cycles of 3-legged cycles. These contour paths had no edge subdivided and consequently we did not subdivide any edge having one between $d_{1}$, $d_{2}$, or $d_{3}$. Hence, there are at least 3 edges subdivided in $G^{1}_{C}$ and by Lemma C.1, there are no bad 2-legged cycles.
We now prove that there are no pairwise intersecting demanding cycle contained in $G_{C}$. Suppose by contradiction that it is the case. Then, we slightly modify graph $G$ so that we have a contradiction to Condition (3) of Lemma C.2. We proceed as follow: we consider any edge not contained in $G_{C}$ and incident to $f_{\operatorname{ext}}$ and we subdivide it twice inserting two vertices $v_{1}$ and $v_{2}$; we subdivide any other edge incident to the same internal face with a vertex $v_{3}$; We add a vertex $v_{4}$ and the edges $v_{1}v_{4}$, $v_{2}v_{4}$, and $v_{3}v_{4}$ so that the graph we obtain is plane. The obtained graph is also plane triconnected cubic and it has a 3-legged cycle formed by three edges $\{v_{1}v_{2},v_{1}v_{4},v_{2}v_{4}\}$ incident to the external face of $G$ and not contained in $G_{C}$. Such 3-legged cycle is a demanding cycle, since no vertex is contained in its interior. Since $I(G)\not=\emptyset$ by hypothesis, we have a contradiction to Condition (3) of Lemma C.2.
@@
See 4.3
Proof C.20.
We prove that we can compute in $O(n)$ time a collection with at most $q=2|D(G)|+|D_{\exp}(G)|+8-\min\{8,2|D_{\operatorname{ext}}(G)|-|D_{\operatorname{short}}(G)|\}$ bends. Our strategy is first creating two copies of $G$, $G^{1}$ and $G^{2}$, and then placing on them overall exactly $q$ dummies. We prove that none of the two graphs has a bad cycle; then, by Lemma 4.1 and since $q\leq\operatorname{tbn}(G)$ by Lemma 4.2, this directly gives us an unbent collection with $\operatorname{tbn}(G)$ bends. Let $f_{\operatorname{ext}}$ be the external face of $G$. We first prove the theorem for a simple case, where there is no 3-legged cycle with edges incident to $f_{\operatorname{ext}}$ (hence, in particular, $D_{\operatorname{ext}}(G)=\emptyset$) and $f_{\operatorname{ext}}$ is incident to at least four edges. If these conditions hold, we say that $G$ is nice.
G is nice. Let $G^{1}$ and $G^{2}$ be two copies of $G$. Observe that in this case, since $G$ is nice, $D_{\operatorname{ext}}(G)=\emptyset$ and consequently $02|D_{\operatorname{ext}}(G)|-|D_{\operatorname{short}}|=$ 0. Hence, by Lemma 4.2, $\operatorname{tbn}\geq q=2|D(G)|+|D_{\exp}(G)|+8$. We show how to place $q$ dummies overall so that $G^{1}$ and $G^{2}$ have no bad cycle.
We compute the maximal 3-legged cycles of $G$ by Lemma C.10. Let $C$ be one of such maximal 3-legged cycles. Consider the subgraphs $G^{1}_{C}$ and $G^{2}_{C}$ of $G^{1}$ and $G^{2}$, respectively. We place dummies along the edges of $G^{1}_{C}$ and $G^{2}_{C}$ by the algorithm whose existence is stated by Lemma C.18. We perform this procedure for every maximal 3-legged cycle. We have that since $G$ is nice and by Lemma C.2(3), the maximal 3-legged cycles of $G$ are disjoint. Hence, by Lemma C.18, we placed exactly $q^{\prime}=2|D(G)|+|D_{\exp}(G)|$ dummies along the edges of $G^{1}$ and of $G^{2}$. Consider the external face of $G$ and select four edges $e_{1}^{1},e_{2}^{1},e_{1}^{2},e_{2}^{2}$ incident to $f_{\operatorname{ext}}$ such that both the pair $e_{1}^{1},e_{2}^{1}$ and the pair $e_{1}^{2},e_{2}^{2}$ are independent pairs of edges (this is always possible, as $G$ is nice and consequently it has at least four edges incident to the external face). We place two dummies along $e_{1}^{1}$ and $e_{2}^{1}$ in $G^{1}$ and two dummies along $e_{2}^{1}$ and $e_{2}^{2}$ in $G^{2}$. The external cycle of both $G^{1}$ and $G^{2}$ is not a bad cycle, as we just placed four dummies in each of them. Also, there is no other bad cycle by Lemma C.1. Observe that this procedure places $q=2|D(G)|+|D_{\exp}(G)|+8$ dummies. The linear time complexity follows from Lemma 4.1, by the fact that when $G$ is nice maximal 3-legged cycles of $G$ are independent (as observed before), and by Lemma C.18.
@@
We first observe that these cases are mutually exclusive. The fact that (i) and (iii) are mutually exclusive directly comes from Lemma C.2(3). Concerning (ii) and the other two, it suffices to prove that when the external cycle is a 3-cycle there is no 3-legged incident to $f_{\operatorname{ext}}$. Suppose by contradiction that $f_{\operatorname{ext}}$ is incident to 3 edges and that there is a 3-legged $C$ having $f_{\operatorname{ext}}$ as leg face. Then two of the edges incident to $f_{\operatorname{ext}}$ are legs of $C$ and at least one of them is an edge of $C$ forming one of its contour paths. Since these are 3 edges already, incident to $f_{\operatorname{ext}}$, it means that all the 3 legs of $C$ are incident to a same vertex. A contradiction (see the definition of $k$-legged cycle).
(i) Intersecting Case. Refer to Figure 17, where $C_{1}$, $C_{2}$ and $C_{3}$ are intersecting demanding cycles and $C_{1}^{t}$, $C_{2}^{t}$, and $C_{3}^{t}$ are their respective twins. We compute all the intersecting demanding cycles by Lemma C.10, computing all the demanding and subtracting from this set the demandings in $D(G)$. We select two intersecting demanding cycles $C_{1}$ and $C_{2}$ and the relative twins $C_{1}^{t}$ and $C_{2}^{t}$. We then consider $C_{1}$ and $C_{1}^{t}$, we select two edges $e\in C_{1}\cap f_{\operatorname{ext}}$ and $e^{\prime}\in C_{1}^{t}\cap f_{\operatorname{ext}}$. We place in $G^{1}$ two dummies in $e$ and two dummies in $e^{\prime}$. By Lemma C.2(2) there is no 3-legged bad cycle that is intersecting in $G^{1}$. Also the external face is now not a bad cycle in $G^{1}$ and by Lemma C.1 there is no bad 2-legged cycle. We do the same considering $C_{2}$, $C_{2}^{t}$, and $G^{2}$. Concerning the internal 3-legged cycles, we handle them with the nice-placement. By Lemma C.2(3) we have $D_{\operatorname{ext}}(G)=\emptyset$. Since $D_{\operatorname{ext}}(G)=\emptyset$, as in the case when $G$ is nice, we have $\operatorname{tbn}(G)\geq q=2|D(G)|+|D_{\exp}(G)|+8$ by Lemma 4.2. We have that with this placement strategy we placed $q^{\prime}=2|D(G)|+|D_{\exp}(G)|$ internal dummies by Lemma C.18 as in the case when $G$ is nice and, considering the $8$ we placed in the external face to handle the intersecting demanding cycles, we placed overall $q=q^{\prime}+8=2|D(G)|+|D_{\exp}(G)|+8$. Since we have no bad cycle, as we handled the intersecting applying Lemma C.2(2) and since the 3-legged are handled by the same procedure used when $G$ is nice, the lemma follows. Concerning the computational time, the fact that this procedure takes linear time comes from Lemma C.10 and since the nice-placement takes linear time (as proved in the case when $G$ is a nice graph).
@@
(ii) 3-Cycle Case. As observed while showing that the three cases we are considering are mutually exclusive, we have that in this case there is no 3-legged cycle having $f_{\operatorname{ext}}$ as leg face and consequently no intersecting cycle and $D_{\operatorname{ext}}(G)=\emptyset$. We now have 3 edges in the external face and 4 dummies to place so that it is not a bad cycle, but here, the situation changes with respect to the case when $G$ is nice, since we cannot directly use Lemma C.1. In fact, since the external face has 3 edges we cannot place the 4 dummies in 3 different edges or in at least 2 edges with 2 dummies each that are independent, as described by Lemma C.1. Hence, here we do not use directly Lemma 4.2, but we rely on the fact that the nice-placement places dummies optimally in $G^{1}\setminus e$ and $G^{2}\setminus e$ and on the fact that the placement of dummies on edges of $f_{\operatorname{ext}}$ has bounded number of possibilities. Namely, there are three possibilities:
@@
We first show how to handle the 1-2-placement. Consider first $G^{1}$. Refer to Figure 18(a). We subdivide $e=uv$ with four dummies $d_{1},...,d_{4}$ so that $d_{1}$ is adjacent to $u$ and $d_{4}$ is adjacent to $v$. Let $C_{\operatorname{bad}}$ be the external cycle of $G^{1}\setminus e$. We have that $C_{\operatorname{bad}}$ is a bad 2-legged cycle and we need to place two dummies on it. We are going to show how to subdivide $C_{\operatorname{bad}}$ optimally. Observe that $C_{\operatorname{bad}}$ has two legs, which are $(u,d_{1})$ and $(v,d_{4})$. Suppose that $e^{\prime}$ and $e^{*}$, which are the two edges different from $e$ incident to $f_{\operatorname{ext}}$, are incident to $u$ and $v$ respectively. Let $e^{\prime}_{\mathrm{int}}$ and $e^{*}_{\mathrm{int}}$ be the other edges incident to $u$ and $v$, which are internal edges of $G$ Let $\Pi$ be a path such that $C_{\operatorname{bad}}=e^{\prime}\cup e^{*}\cup e^{*}_{\mathrm{int}}\cup e^{\prime}_{\mathrm{int}}\cup\Pi$ and let $f_{e}$ be the face incident to $e$ and internal in $G$.
We now use the nice-placement in $G\setminus e$. We just placed $q^{\prime}=2|D(G)|+|D_{\exp}(G)|$. By Lemma C.8 the interesting contour paths and, consequently, the legs of the demanding cycles incident to $f_{e}$ lie along $\Pi$. Hence, if such demanding cycles exists, nice-placement subdivided $\Pi$ already by Lemma C.18(2). If $C_{\operatorname{bad}}$ has 2 dummies in $\Pi$, we do nothing, and the cost is optimum as we did not add dummies with respect to the nice-placement. Otherwise, suppose that that $C_{\operatorname{bad}}$ does not have 2 dummies.
@@
We now consider $G^{2}$. Refer to Figure 18(b). Let $w$ be the vertex of $G$ incident to $e^{\prime}$ and $e^{*}$. We have to consider the degenerate 3-legged cycle $C_{\operatorname{bad}}^{\prime}$ of $G$ whose three legs are incident to $w$, since two of its legs are incident to dummies. In this case we have to place just one dummy $d^{\prime}$. As before, we can observe that either the nice-placement placed such dummy in one of its contour paths $\Pi^{\prime}$ and $\Pi^{*}$ (sharing a face with $e^{\prime}$ and $e^{*}$, respectively, and defined similarly as $\Pi$) previously, or we can place it in any edge of such contour paths. We have that we did not place a dummy in a same edge in $G^{1}$ and $G^{2}$. In fact, $\Pi^{\prime}$ and $\Pi^{*}$ are disjoint from $\Pi$.
@@
(iii) External-Demanding Case: $D_{\operatorname{ext}}(G)\not=\emptyset$. As there is no intersecting cycle, since the cases we are considering are mutually exclusive, we have that the maximal 3-legged cycles are independent. Hence, we can use the nice-placement so that now every 3-legged cycle is not a bad cycle. We placed exactly $q^{\prime}=2|D(G)|+|D_{\exp}(G)|$ dummies. We have now to see what happens in the external face $f_{\operatorname{ext}}$. By Lemma C.18(2) we placed a dummy in each of $G^{1}$ $G^{2}$ for every cycle in $D_{\operatorname{ext}}(G)\setminus D_{\operatorname{short}}(G)$ and just one dummy (either in $G^{1}$ or $G^{2}$) for the cycles in $D_{\operatorname{short}}(G)$. The remaining dummies so that the external face has at least four vertices can be placed in other edges in $G^{1}$ and $G^{2}$. We have that by Lemma C.18 and by Lemma 4.2 we placed $\operatorname{tbn}(G)$ dummies. Also, by Lemma C.2 and by Lemma C.18 there is no bad cycle.
@@
Corollary C.21.
Let $G$ be a planar triconnected cubic with $n$ vertices. There exists an $O(n^{2})$-time algorithm that computes an unbent collection of $G$ with $\operatorname{tbn}(G)$ bends. Also, the collection has size two which is optimal.